# Two-dimensional fluids via matrix hydrodynamics

### v2

@@ 2. Background: geometry of the 2-D Euler equations
In addition to Casimirs, the transport property implies weak (distributional), finite-dimensional co-adjoint orbits given by $n$ interacting point vortices
$\omega=\sum_{k=1}^{n}\Gamma_{k}\delta_{x_{k}}.$
These were found by Helmholtz [14], and analyzed by Kirchhoff [23] and Kelvin [20]. Much later, Onsager [33] used them is his theory of statistical hydrodynamics. The existence of weak solutions shows that the restriction to the smooth dual is not always legitimate.
Instead of $\omega_{t}$ as dynamic variable, we can use the transport map $\Phi_{t}\in\operatorname{Diff}_{\mu}(\mathbb{S}^{2})$ via the equation
$\dot{\Phi}_{t}=X_{\psi}\circ\Phi_{t},\quad\Delta\psi=\omega_{0}\circ\Phi_{t}^{-1}.$ (6)
This is the transport map formulation of equation (5). Notice that $\omega$ is not required to be differentiable here, which is the key to Yudovich’s [39] global well-posedness results in $L^{\infty}$.
What happens with the Hamiltonian structure when we switch to the transport formulation in (6)? Interestingly, the group $\operatorname{Diff}_{\mu}(\mathbb{S}^{2})$ inherits the symplectic structure of $\mathbb{S}^{2}$. Indeed, as an infinite-dimensional Lie group, the tangent bundle $TG$ is trivializable as $G\times\mathfrak{g}$ via right translation. The symplectic structure is then
$T_{\Phi}G\times T_{\Phi}G\simeq\mathfrak{g}\times\mathfrak{g}\ni(\psi,\xi)\mapsto\int_{\mathbb{S}^{2}}\{\psi,\xi\}\,\mu.$
With this symplectic structure, equation (6) is Hamilton’s equation on $G=\operatorname{Diff}_{\mu}(\mathbb{S}^{2})$ for the Hamiltonian $\bar{H}(\Phi)=H(\omega_{0}\circ\Phi^{-1})$.
@@
We have thus summarized the complete geometric description of the 2-D Euler equations on $\mathbb{S}^{2}$. Indeed, the structures presented – the Hamiltonian formulation, the co-adjoint orbits, the conservation of Casimirs, and the transport formulation – are all important in the various mathematical theories for long-time behavior (cf. surveys by Drivas and Elgindi [9] and Khesin, Misiołek, and Shnirelman [21]).
@@ 3. Quantization and the Euler–Zeitlin equation
    such that, in some sense, the Poisson bracket is approximated by the matrix commutator:
    $T_{N}(\{\psi,\omega\})\to\frac{1}{\hbar}[T_{N}\psi,T_{N}\omega]\quad\text{as}\quad N\to\infty,$
@@
    The projection maps $T_{N}$ should take the unit function $x\mapsto 1$ to the imaginary identity matrix $iI\in\mathfrak{u}(N)$. In particular, $T_{N}$ descends to a map between the quotient Lie algebras $C^{\infty}(\mathbb{S}^{2})/\mathbb{R}$ and $\mathfrak{pu}(N)\equiv\mathfrak{u}(N)/i\mathbb{R}I$ [^1] We refrain from identifying the Lie algebra $\mathfrak{pu}(N)$ with $\mathfrak{su}(N)$, since its Lie group $\mathrm{PU}(N)$ is different from $\mathrm{SU}(N)$.
@@
    $\operatorname{ad}^{*}_{P}W=-\frac{1}{\hbar}[P,W].$
@@
    $\dot{W}+\frac{1}{\hbar}[P,W]=0,\qquad\Delta_{N}P=W.$ (7)
@@ 3.1. Quantization via representation
To build the quantization maps $T_{N}$, it is natural to begin from the sphere’s symmetry: the Lie group $\operatorname{SO}(3)$. The connection to the space of functions $C^{\infty}(\mathbb{S}^{2})$ is established via its Lie algebra $\mathfrak{so}(3)$ as follows. Let $x_{1},x_{2},x_{3}\in C^{\infty}(\mathbb{S}^{2})$ be a choice of Euclidean coordinate functions for the embedded sphere $\mathbb{S}^{2}\subset\mathbb{R}^{3}$. These functions generate $\mathfrak{so}(3)$ as a Lie sub-algebra of the Poisson algebra $\mathfrak{g}=(C^{\infty}(\mathbb{S}^{2}),\{\cdot,\cdot\})$. But we should think of it differently. Namely, the basis elements $\mathbf{e}_{1},\mathbf{e}_{2},\mathbf{e}_{3}\in\mathbb{R}^{3}\simeq\mathfrak{so}(3)$ are associated with first order differential operators $\mathcal{X}_{i}$ on $C^{\infty}(\mathbb{S}^{2})$ given by
$\mathcal{X}_{\alpha}\omega=\{x_{\alpha},\omega\}.$
@@
Then $\mathcal{X}_{1},\mathcal{X}_{2},\mathcal{X}_{3}$ provides a unitary representation of $\mathfrak{so}(3)$ on $C^{\infty}(\mathbb{S}^{2})$ compatible with the Poisson algebra structure. Indeed, the Jacobi identity yields $\mathcal{X}_{\alpha}\mathcal{X}_{\beta}-\mathcal{X}_{\beta}\mathcal{X}_{\alpha}=\{\{x_{\alpha},x_{\beta}\},\cdot\}$ and each $\mathcal{X}_{\alpha}$ is skew self-adjoint with respect to the $L^{2}$ inner product. We now apply the powerful machinery of Schur, Cartan, Weyl, and others (cf. [25]) for representations of compact groups and the connection to harmonic analysis. First, the generators $\mathcal{X}_{1},\mathcal{X}_{2},\mathcal{X}_{3}$ can be viewed as the first order elements of the universal enveloping algebra $U(\mathfrak{so}(3))$. Higher order elements correspond to higher order differential operators. In particular, the Killing form on $\mathfrak{so}(3)$ corresponds to the second order Casimir element $\mathcal{C}=-\sum_{\alpha}\mathcal{X}_{\alpha}^{2}$, which in this case is the Laplacian $\Delta$ on $\mathbb{S}^{2}$. The Casimir element commutes with each $\mathcal{X}_{\alpha}$. In particular, if $V\subset C^{\infty}(\mathbb{S}^{2})$ is an eigenspace of $\mathcal{C}=\Delta$ with eigenvalue $\lambda$, then $\mathcal{X}_{1},\mathcal{X}_{2},\mathcal{X}_{3}$ provide a sub-representation of $\mathfrak{so}(3)$ on $V$. Indeed, $\mathcal{C}[\mathcal{X}_{\alpha},\mathcal{X}_{\beta}]V=[\mathcal{X}_{\alpha},\mathcal{X}_{\beta}]\mathcal{C}V=\lambda[\mathcal{X}_{\alpha},\mathcal{X}_{\beta}]V$, so $[\mathcal{X}_{\alpha},\mathcal{X}_{\beta}]V$ must belong to the eigenspace of $\lambda$, i.e., $[\mathcal{X}_{\alpha},\mathcal{X}_{\beta}]V\subset V$. One can further prove that this representation is irreducible. In fact, the eigenspaces $V_{\ell}$ for $\lambda_{\ell}=-\ell(1+\ell),\ell=0,1,\ldots$ correspond to all the irreducible infinitesimal representations of $\mathrm{SO}(3)$ (i.e., the odd-dimensional representations of $\mathfrak{so}(3))$. Furthermore, $V_{\ell}$ has dimension $12\ell+$ 1 and comes with a canonical basis $Y_{\ell m}$ (the spherical harmonics) corresponding to the weights $m=-\ell,\ldots,\ell$ of the representation.
@@
For $N$ even or odd, corresponding to $\ell^{\prime}=(N-1)/2$, begin with an irreducible, unitary representation $\pi\colon\mathfrak{so}(3)\to\mathfrak{u}(N)$. Let $S_{\alpha}=\pi(\mathbf{e}_{\alpha})$, chosen so that $S_{3}$ is diagonal with sorted, equidistant eigenvalues $-\mathrm{i}\ell^{\prime},\ldots,\mathrm{i}\ell^{\prime}$ corresponding to the weights of the representation. Since the representation is unitary, $S_{1},S_{2},S_{3}$ share the same spectrum.
The Casimir element restricted to the representation is just multiplication by $-\ell^{\prime}(1+\ell^{\prime})$ and is therefore given by the diagonal matrix $C=-\sum_{\alpha}S_{\alpha}^{2}=-\ell^{\prime}(1+\ell^{\prime})I$. We now introduce a scaled representation $X_{\alpha}=\hbar S_{\alpha}$ in such a way that the matrix $X_{\alpha}$ correspond to the function $x_{\alpha}$ on $\mathbb{S}^{2}$. Thus, to achieve $\big{(}\sum_{\alpha}X_{\alpha}^{2}\big{)}^{1/2}=\mathrm{i}I$, corresponding to $\big{(}\sum_{\alpha}x_{\alpha}^{2}\big{)}^{1/2}=1$, we take $\hbar=1/\sqrt{\ell^{\prime}(1+\ell^{\prime})}=2/\sqrt{N^{2}-1}\eqqcolon\hbar_{N}$.
@@
Being the Casimir element, any eigenvalue is of the form $\lambda=-\ell(1+\ell)$ for some $\ell\in\mathbb{N}/2$ and the corresponding eigenspace $V_{\ell}$ has dimension $12\ell+$ 1. From the Peter–Weyl theorem, the representation on $\mathfrak{u}(N)$ decomposes into irreducible, orthogonal components $\mathfrak{u}(N)\simeq V_{\ell_{0}}\oplus\cdots\oplus V_{\ell_{n-1}}$. But which $\ell_{k}$ are included?
Since $\Delta_{N}\mathrm{i}I=0$ and $\Delta_{N}X_{\alpha}=2X_{\alpha}$, we have $\ell_{0}=0$ and $\ell_{1}=1$. More generally and as desired, $\ell_{k}=k$ and $n=N$:
@@
The components $V_{\ell m}$, corresponding to the weights ±$\pm 2m$, are mapped to the ±mplus-or-$\pm m$th diagonals of the matrix in $\mathfrak{u}(N)$.
@@
The decomposition in Lemma 1 enables explicit construction of the projection map $T_{N}\colon\mathfrak{g}\to\mathfrak{u}(N)$. Indeed, the ±mplus-or-$\pm m$th diagonals of $W\in\mathfrak{u}(N)$ are spanned by $V_{:m}\coloneqq V_{mm}\oplus\cdots\oplus V_{(N-1)m}\simeq\mathbb{C}^{N-m}$. In turn, $V_{:m}$ corresponds to the span of the spherical harmonic functions $Y_{mm},\ldots,Y_{(N-1)m}$. Since $\Delta_{N}V_{:m}\subseteq V_{:m}$, we obtain the ±mplus-or-$\pm m$ diagonals by identifying the eigenvector of $\Delta_{N}|_{V_{:m}}$ with eigenvalue $-(m+k)(1+m+k)$ with the spherical harmonic function $Y_{(m+k)m}$. Specifically, $\Delta_{N}|_{V_{:m}}\colon\mathbb{C}^{N-m}\to\mathbb{C}^{N-m}$, as an operator acting on the upper $m$th diagonal of $W$, is a tridiagonal, symmetric matrix. The spectral decomposition of $\Delta_{N}|_{V_{:m}}$ can hence be computed by efficient numerical methods of complexity $\mathcal{O}(N)$. This way, a function $\omega\in\mathfrak{g}$ is projected to the matrix $W=T_{N}\omega$ from its spherical harmonics coefficients $(\omega_{\ell m})_{\ell\leq N-1}$ in only $\mathcal{O}(N^{2})$ numerical operations; see Cifani et al. [5] for details. Notice that the $L^{2}$ dual $T_{N}^{*}\colon\mathfrak{u}(N)\to\mathfrak{g}$ is a right inverse of $T_{N}$.
@@
Zeitlin’s original model on the flat torus does not correspond to a representation of the symmetry group, which implies more of Gibb’s phenomenon than in the sphere model.
[…]
@@
$\lVert\omega\rVert_{L^{\infty}}-\alpha\hbar\lVert\omega\rVert_{C^{2}}\leq\lVert T_{N}\omega\rVert_{\infty}\leq\lVert\omega\rVert_{L^{\infty}},$
where $\lVert W\rVert_{\infty}=\sup_{\mathbf{v}}\frac{\lvert W\mathbf{v}\rvert}{\lvert\mathbf{v}\rvert}$. Further, there exists $\beta>0$ so for all $\omega,\psi\in C^{3}(\mathbb{S}^{2})$
@@
Remark 3.
Since $C^{2}(\mathbb{S}^{2})$ is dense in $L^{\infty}(\mathbb{S}^{2})$, it follows from the first inequality in Theorem 3 that $T_{N}$ can be extended to a bounded operator $L^{\infty}(\mathbb{S}^{2})\to\mathfrak{su}(N)$.
@@
See appendix. ∎
@@
The result [^2] See also Theorem 16 of the appendix for a more detailed statement. in Theorem 4 is analogous to the $L^{2}$ convergence result of Gallagher [13] for Zeitlin’s original model on the flat torus $\mathbb{T}^{2}$. But there are fundamental differences in the techniques: Gallagher’s result rely on the algebraic multiplication properties of the Fourier basis in $\mathbb{T}^{2}$, which are not available on $\mathbb{S}^{2}$.

### v3

@@ 2. Background: geometry of the 2-D Euler equations
The transport property furthermore implies weak (distributional), finite-dimensional coadjoint orbits given by $n$ interacting point vortices
$\omega=\sum_{k=1}^{n}\Gamma_{k}\delta_{x_{k}}.$ (7)
These were found by Helmholtz [25], and analyzed by Kirchhoff [35] and Kelvin [31]. Much later, Onsager [48] used them is his theory of statistical hydrodynamics. Let us see how they fit into the geometric framework of Lie-Poisson dynamics, following Marsden and Weinstein [43]. In general, if $G$ acts on a symplectic manifold $(M,\nu)$ by symplectomorphisms then there is a corresponding infinitesimal action vector field $\xi_{M}\in\mathfrak{X}_{\nu}(M)$ of $\xi\in\mathfrak{g}$. Suppose that $\xi_{M}$ is Hamiltonian, i.e., there is a function $H_{\xi}\in C_{0}^{\infty}(M)$ such that $\iota_{\xi_{M}}\nu=\mathrm{d}H_{\xi}$. Since $H_{\xi}$ is linear in $\xi$, we have that
$H_{\xi}(m)=\langle p(m),\xi\rangle$
for some function $p\colon M\to\mathfrak{g}^{*}$ called the momentum map. Usually (but not always) the momentum map is equivariant, i.e., $p(g\cdot m)=\mathrm{Ad}^{*}_{g}(p(m))$. If, in addition, the action of $G$ on $M$ is transitive then the image $p(M)\subset\mathfrak{g}^{*}$ is a (generally weak) coadjoint orbit.
[…]
Since the action is transitive, we expect corresponding finite dimensional coadjoint orbits. Indeed, $\xi_{M}(x_{1},\ldots,x_{n})=\big{(}\xi(x_{1}),\ldots,\xi(x_{n})\big{)}$ so
$\iota_{\xi_{M}}\nu=\sum_{k}\Gamma_{k}\iota_{\xi(x_{k})}\mu=\sum_{k}\Gamma_{k}\mathrm{d}\psi(x_{k})=\mathrm{d}\sum_{k}\Gamma_{k}\psi(x_{k})=\mathrm{d}\int_{\mathbb{S}^{2}}\underbrace{\left(\sum_{k}\Gamma_{k}\delta_{x_{k}}\right)}_{\omega=p(x_{1},\ldots,x_{n})}\psi\mu,$
where $\xi=X_{\psi}$. Thus, the associated finite dimensional (weak) coadjoint orbit is given by the point vortices (7). These orbits and the corresponding weak solutions show that the restriction to the smooth dual is not always legitimate.
To make the analogy with continuous vorticity functions, one should think of phase space for point vortices as parameterized by $(x_{1},\ldots,x_{n},\Gamma_{1},\ldots,\Gamma_{n})$. The positions $x_{k}$ correspond to the level sets of a smooth vorticity function, and the strengths $\Gamma_{k}$ to the values associated with the level sets. Just like the function values of the initial conditions $\omega_{0}\in C^{\infty}(\mathbb{S}^{2})$ are preserved (since $\omega_{t}$ is transported), the strengths $\Gamma_{k}$ are conserved. The dynamics for point vortices take place in the map $(x_{1}(0),\ldots,x_{n}(0))\mapsto(x_{1}(t),\ldots,x_{n}(t))$ which moves the strengths around. The coordinates $(x_{1},\ldots,x_{n},\Gamma_{1},\ldots,\Gamma_{n})$ for point vortices have a direct analog for smooth coadjoint orbits. Indeed, instead of $\omega_{t}$ as dynamic variable, we can use the map $\Phi_{t}\in\operatorname{Diff}_{\mu}(\mathbb{S}^{2})$ together with the initial conditions $\omega_{0}$ via the equation
$\dot{\Phi}_{t}=X_{\psi}\circ\Phi_{t},\quad-\Delta\psi=\omega_{0}\circ\Phi_{t}^{-1}.$ (8)
This is the transport map formulation of equation (6). It is the basis for Yudovich’s [59] global well-posedness results in $L^{\infty}$, which shows convergence of the fixed point iteration obtained by leap-frogging between the equations (8) for $\Phi$ (Picard iterations for ODE) and for $\psi$ (stationary PDE). Geometrically, the transport formulation captures that the symplectic leaves are generated by the action of the Lie group.
@@
We have thus summarized the geometric description of the 2-D Euler equations on $\mathbb{S}^{2}$. Indeed, the structures presented – the Hamiltonian formulation, the coadjoint orbits, the conservation of Casimirs, and the transport formulation – are all important in the various mathematical theories for long-time behavior (cf. surveys by Drivas and Elgindi [18] and Khesin, Misiołek, and Shnirelman [32]).
Remark 1.
Arnold’s [2] geometric description is valid for Euler’s equations on Riemannian manifolds $M$ of arbitrary dimension, by taking the group $G$ to be volume preserving diffeomorphisms $\mathrm{Diff}_{\mathrm{vol}}(M)$ with the Lie algebra $\mathfrak{g}$ of divergence free vector fields. However, this setting fails to capture some geometric qualities of 2-D Euler. In particular, the action of $\mathrm{Diff}_{\mathrm{vol}}(M)$ on $M$ is not symplectic if $\operatorname{dim}M>2$, nor is the action on $C^{\infty}(M)$ a Poisson map, which, for example, imply that singular and smooth coadjoint orbits cannot be given as point vortices and vorticity functions. Instead, the geometrically natural generalization of 2-D Euler is the group of (Hamiltonian) symplectomorphisms of a symplectic manifold equipped with a Riemannian metric (e.g., a Kähler manifold). This setting captures all the geometry of the 2-D Euler equations and therefore many of the analysis results (e.g., global solutions [19]).
@@ 3. Quantization and the Euler–Zeitlin equation
    such that the Poisson bracket is approximated by the matrix commutator in the spectral norm $\lVert\cdot\rVert_{\infty}$:
    $\lVert T_{N}(\{\psi,\omega\})-\frac{1}{\hbar}[T_{N}\psi,T_{N}\omega]\rVert_{\infty}\to 0\quad\text{as}\quad N\to\infty,$
@@
    The projection maps $T_{N}$ should take the unit function $x\mapsto 1$ to the imaginary identity matrix $iI\in\mathfrak{u}(N)$. In particular, $T_{N}$ descends to a map between the quotient Lie algebras $C^{\infty}(\mathbb{S}^{2})/\mathbb{R}$ and $\mathfrak{pu}(N)\equiv\mathfrak{u}(N)/i\mathbb{R}I$. [^2] We refrain from identifying the Lie algebra $\mathfrak{pu}(N)$ with $\mathfrak{su}(N)$, since its Lie group $\mathrm{PU}(N)$ is different from $\mathrm{SU}(N)$.
@@
    $\operatorname{ad}^{*}_{P}W=\frac{1}{\hbar}[W,P].$
@@
    $\dot{W}+\frac{1}{\hbar}[W,P]=0,\qquad-\Delta_{N}P=W.$ (9)
@@ 3.1. Quantization via representation
To build the quantization maps $T_{N}$, it is natural to begin from the sphere’s symmetry: the Lie group $\operatorname{SO}(3)$. The connection to the space of functions $C^{\infty}(\mathbb{S}^{2})$ is established via its Lie algebra $\mathfrak{so}(3)$ as follows. Let $x_{1},x_{2},x_{3}\in C^{\infty}(\mathbb{S}^{2})$ be a choice of Euclidean coordinate functions for the embedded sphere $\mathbb{S}^{2}\subset\mathbb{R}^{3}$. These functions generate $\mathfrak{so}(3)$ as a Lie subalgebra of the Poisson algebra $\mathfrak{g}=(C^{\infty}(\mathbb{S}^{2}),\{\cdot,\cdot\})$. But we should think of it differently. Namely, the basis elements $\mathbf{e}_{1},\mathbf{e}_{2},\mathbf{e}_{3}\in\mathbb{R}^{3}\simeq\mathfrak{so}(3)$ are associated with first order differential operators $\mathcal{X}_{\alpha}$ on $C^{\infty}(\mathbb{S}^{2})$ given by
$\mathcal{X}_{\alpha}\omega=\{x_{\alpha},\omega\},\qquad\text{for}\;\alpha=1,2,3\,.$
@@
Then $\mathcal{X}_{1},\mathcal{X}_{2},\mathcal{X}_{3}$ provides a unitary representation of $\mathfrak{so}(3)$ on $C^{\infty}(\mathbb{S}^{2})$ compatible with the Poisson algebra structure. Indeed, the Jacobi identity yields $\mathcal{X}_{\alpha}\mathcal{X}_{\beta}-\mathcal{X}_{\beta}\mathcal{X}_{\alpha}=\{\{x_{\alpha},x_{\beta}\},\cdot\}$ and each $\mathcal{X}_{\alpha}$ is skew self-adjoint with respect to the $L^{2}$ inner product. We now apply the powerful machinery of Schur, Cartan, Weyl, and others (cf. [38]) for representations of compact groups and the connection to harmonic analysis. First, the generators $\mathcal{X}_{1},\mathcal{X}_{2},\mathcal{X}_{3}$ can be viewed as the first order elements of the universal enveloping algebra $U(\mathfrak{so}(3))$. Higher order elements correspond to higher order differential operators. In particular, the Killing form on $\mathfrak{so}(3)$ corresponds to the second order Casimir element $\mathcal{C}=-\sum_{\alpha}\mathcal{X}_{\alpha}^{2}$, which in this case is the Laplacian $\Delta$ on $\mathbb{S}^{2}$. The Casimir element commutes with each $\mathcal{X}_{\alpha}$. In particular, if $V\subset C^{\infty}(\mathbb{S}^{2})$ is an eigenspace of $\mathcal{C}=\Delta$ with eigenvalue $\lambda$, then $\mathcal{X}_{1},\mathcal{X}_{2},\mathcal{X}_{3}$ provide a sub-representation of $\mathfrak{so}(3)$ on $V$. Indeed, $\mathcal{C}[\mathcal{X}_{\alpha},\mathcal{X}_{\beta}]V=[\mathcal{X}_{\alpha},\mathcal{X}_{\beta}]\mathcal{C}V=\lambda[\mathcal{X}_{\alpha},\mathcal{X}_{\beta}]V$, so $\mathcal{X}_{\alpha},\mathcal{X}_{\beta}]V$ must belong to the eigenspace of $\lambda$, i.e., $\mathcal{X}_{\alpha},\mathcal{X}_{\beta}]V\subset V$. One can further prove that this representation is irreducible. In fact, the eigenspaces $V_{\ell}$ for $\lambda_{\ell}=-\ell(1+\ell),\ell=0,1,\ldots$ correspond to all the irreducible infinitesimal representations of $\mathrm{SO}(3)$ (i.e., the odd-dimensional representations of $\mathfrak{so}(3))$. Furthermore, $V_{\ell}$ has dimension $2\ell+1$ and comes with a canonical basis $Y_{\ell m}$ (the spherical harmonics) corresponding to the weights $m=-\ell,\ldots,\ell$ of the representation.
@@
For $N$ even or odd, corresponding to $\ell^{\prime}=(N-1)/2$, begin with an irreducible, unitary representation $\pi\colon\mathfrak{so}(3)\to\mathfrak{u}(N)$. Let $S_{\alpha}=\pi(\mathbf{e}_{\alpha})$, chosen so that $S_{3}$ is diagonal with sorted, equidistant eigenvalues $\mathrm{i}\ell^{\prime},\ldots,\mathrm{i}\ell^{\prime}$ corresponding to the weights of the representation. Since the representation is unitary, $S_{1},S_{2},S_{3}$ share the same spectrum. Explicitly, $S_{1}$ and $S_{2}$ are the tridiagonal matrices given by
$S_{1}=\frac{\mathrm{i}}{2}\begin{pmatrix}0&&&\\ &\ddots&f(w)&\\ &f(w)&\ddots&\\ &&&0\\ \end{pmatrix}\quad\text{and}\quad S_{2}=\frac{\mathrm{1}}{2}\begin{pmatrix}0&&&\\ &\ddots&f(w)&\\ &-f(w)&\ddots&\\ &&&0\\ \end{pmatrix}$
for $w=-\ell^{\prime},-\ell^{\prime}+1,\ldots,\ell^{\prime}-1$ and $f(w)=\sqrt{\ell^{\prime}(\ell^{\prime}+1)-w(w+1)}$.
The Casimir element restricted to the representation is just multiplication by $\ell^{\prime}(1+\ell^{\prime})$ and is therefore given by the diagonal matrix $C=-\sum_{\alpha}S_{\alpha}^{2}=-\ell^{\prime}(1+\ell^{\prime})I$. We now introduce a scaled representation $X_{\alpha}=\hbar S_{\alpha}$ in such a way that the matrix $X_{\alpha}$ correspond to the function $x_{\alpha}$ on $\mathbb{S}^{2}$. Thus, to achieve $\big{(}\sum_{\alpha}X_{\alpha}^{2}\big{)}^{1/2}=\mathrm{i}I$, corresponding to $\big{(}\sum_{\alpha}x_{\alpha}^{2}\big{)}^{1/2}=1$, we take $\hbar=1/\sqrt{\ell^{\prime}(1+\ell^{\prime})}=2/\sqrt{N^{2}-1}\eqqcolon\hbar_{N}$.
@@
Being the Casimir element, any eigenvalue is of the form $\lambda=-\ell(1+\ell)$ for some $\ell\in\mathbb{N}/2$ and the corresponding eigenspace $V_{\ell}$ has dimension $2\ell+1$. From the Peter–Weyl theorem, the representation on $\mathfrak{u}(N)$ decomposes into irreducible, orthogonal components $\mathfrak{u}(N)\simeq V_{\ell_{0}}\oplus\cdots\oplus V_{\ell_{n-1}}$. But which $\ell_{k}$ are included?
Since $\Delta_{N}\mathrm{i}I=0$ and $\Delta_{N}X_{\alpha}=2X_{\alpha}$, we have $\ell_{0}=0$ and $\ell_{1}=1$. From general theory, and as desired, $\ell_{k}=k$ and $n=N$:
@@
The components $V_{\ell m}$, corresponding to the weights $\pm 2m$, are mapped to the $\pm m$th diagonals of the matrix in $\mathfrak{u}(N)$.
@@
The result is well-known and follows from general theory about multiplicities in the Peter-Weyl decomposition, see for example Kirillov [36, ch. 4, particularly sec. 4.9]. We give here a direct, less general proof.
@@
The decomposition in Lemma 1 enables explicit construction of the projection map $T_{N}\colon\mathfrak{g}\to\mathfrak{u}(N)$. Indeed, the $\pm m$th diagonals of $W\in\mathfrak{u}(N)$ are spanned by $V_{:m}\coloneqq V_{mm}\oplus\cdots\oplus V_{(N-1)m}\simeq\mathbb{C}^{N-m}$. In turn, $V_{:m}$ corresponds to the span of the spherical harmonic functions $Y_{mm},\ldots,Y_{(N-1)m}$. Since $\Delta_{N}V_{:m}\subseteq V_{:m}$, we obtain the $\pm m$ diagonals by identifying the eigenvector of $\Delta_{N}|_{V_{:m}}$ with eigenvalue (m+k)(1+m+k)-(m+k)(1+m+k) with the spherical harmonic function $Y_{(m+k)m}$. Specifically, $\Delta_{N}|_{V_{:m}}\colon\mathbb{C}^{N-m}\to\mathbb{C}^{N-m}$, as an operator acting on the upper $m$th diagonal of $W$, is a tridiagonal, symmetric matrix. The spectral decomposition of $\Delta_{N}|_{V_{:m}}$ can hence be computed by efficient numerical methods of complexity $\mathcal{O}(N)$. This way, a function $\omega\in\mathfrak{g}$ is projected to the matrix $W=T_{N}\omega$ from its spherical harmonics coefficients $(\omega_{\ell m})_{\ell\leq N-1}$ in only $\mathcal{O}(N^{2})$ numerical operations; see Cifani et al. [11] for details. Notice that the $L^{2}$ adjoint $T_{N}^{*}\colon\mathfrak{u}(N)\to\mathfrak{g}$ is a right inverse of $T_{N}$.
@@
Zeitlin’s original model on the flat torus $\mathbb{T}^{2}$ does not preserve the translational symmetry generated by the action of the Abelian group $A=\mathbb{R}^{2}$ on $\mathbb{R}^{2}/\mathbb{Z}^{2}\simeq\mathbb{T}^{2}$. The problem is that there is no corresponding action of $A$ on the finite dimensional Lie algebra $\mathfrak{g}=\mathfrak{su}(N)$. Indeed, whereas there are actions of $\mathbb{R}$ on $\mathfrak{su}(N)$ corresponding to $(a_{1},0)\in A$ and $(0,a_{2})\in A$, these actions fail to commute. The numerical consequence is that translation-like motion in Zeitlin’s model on $\mathbb{T}$ give rise to artifacts and the Gibbs phenomenon.
The remaining step (1) above, relating the Poisson bracket $\{\cdot,\cdot\}$ to the scaled commutator $\frac{1}{\hbar}[\cdot,\cdot]$, is at the heart of quantization theory. The following is a specialization to our setting of a result by Charles and Polterovich [8].
@@
$\lVert\omega\rVert_{L^{\infty}}-\alpha\hbar\lVert\omega\rVert_{C^{2}}\leq\sqrt{\frac{N-1}{N+1}}\lVert T_{N}\omega\rVert_{\infty}\leq\lVert\omega\rVert_{L^{\infty}},$
where $\lVert W\rVert_{\infty}=\sup_{\mathbf{u}}\frac{\lvert W\mathbf{u}\rvert}{\lvert\mathbf{u}\rvert}$. Further, there exists $\beta>0$ so for all $\omega,\psi\in C^{3}(\mathbb{S}^{2})$
@@
Proof.
The result was proved for Berezin–Toeplitz operators between holomorphic sections of certain complex holomorphic vector bundles appearing in geometric quantization. Up to scaling, the Berezin–Toeplitz quantization is equivalent to the quantization $T_{N}$ discussed here, as explained in Appendix B below. ∎
@@
See Appendix A, where we also give a more detailed formulation of the theorem (see Theorem 19), including convergence rates. ∎
@@
The result in Theorem 4 (and Theorem 19) is analogous to the $L^{2}$ convergence result of Gallagher [23] for Zeitlin’s original model on the flat torus $\mathbb{T}^{2}$. But there are technical differences in the proof: Gallagher’s result rely on the algebraic multiplication properties of the Fourier basis in $\mathbb{T}^{2}$, which are not available on $\mathbb{S}^{2}$.
In summary, we have a correspondence between functions and matrices that allows transcription of ideas and results from 2-D hydrodynamics to matrix theory. The dictionary in Table 1 contains some direct examples.
2-D hydrodynamics matrix hydrodynamics
Lie group $\Phi\in\mathrm{Diff}_{\mu}(\mathbb{S}^{2})$ $F\in\mathrm{PU}(N)$
Lie algebra $\psi\in C^{\infty}(\mathbb{S}^{2})/\mathbb{R},\;\{\cdot,\cdot\}$ $P\in\mathfrak{pu}(N),\;\frac{1}{\hbar}[\cdot,\cdot]$
Phase space $\omega\in C^{\infty}_{0}(\mathbb{S}^{2})$ $W\in\mathfrak{su}(N)$
$\mathrm{Ad}^{*}$ $\omega\circ\Phi^{-1}$ $FWF^{-1}$
$\mathrm{ad}^{*}$ $\{\omega,\psi\}$ $\frac{1}{\hbar}[W,P]$
Strong norm $\lVert\omega\rVert_{L^{\infty}}=\operatorname{ess~sup}\omega$ $\lVert W\rVert_{\infty}=\sup_{\lvert\boldsymbol{v}\rvert=1}\lvert W\boldsymbol{v}\rvert$
Inner product $\langle\omega,\omega\rangle_{L^{2}}=\int_{\mathbb{S}^{2}}\omega^{2}\mu$ $\langle W,W\rangle_{2}=-\frac{4\pi}{N}\operatorname{tr}(W^{2})$
Casimirs $C_{f}(\omega)=\int_{\mathbb{S}^{2}}f(\omega)\mu$ $C_{f}^{N}(W)=\frac{4\pi}{N}\operatorname{tr}(f(-\mathrm{i}W))$
Laplacian $\Delta\psi=\sum_{\alpha=1}^{3}\{x_{\alpha},\{x_{\alpha},\psi\}\}$ $\Delta_{N}P=\frac{1}{\hbar^{2}}\sum_{\alpha=1}^{3}[X_{\alpha},[X_{\alpha},P]]$
Hamiltonian $H(\omega)=-\frac{1}{2}\langle\omega,\Delta^{-1}\omega\rangle_{L^{2}}$ $H_{N}(W)=\frac{2\pi}{N}\operatorname{tr}(W\Delta_{N}^{-1}W)$
Table 1. Dictionary between 2-D and matrix hydrodynamics.